Olmo, Inc., manufactures and sells two products: Product K0 and Product H9. The annual production and sales of Product of K0 is 600 units and of Product H9 is 600 units. The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity:Olmo, Inc., manufactures and sells two products: Product KO and Product H9. The annual production and sales of Product of KO is 600 units and of Product H9 is 600 units. The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity: Activity Estimated Ex Activit Activity Cost Pools Measures Overhead Cost Product Ko Product H9 Tota.l 7,200 900 6,500 Labor-related Production orders Order size $ 550,708 53,719 836,316 $1,440,743 4,800 700 2,900 2,400 200 3,600 DLHS orders MHS The overhead applied to each unit of Product H9 under activity-based costing is closest to: (Round your intermediate calculations to 2 decimal places.) Multiple Choice S1,09782 per unit S77196 per unit $600.31 per unit S305.96 per unit
Answer #1

Overhead cost/activity=Total overhead/Total activity

Hence overhead cost per:

Direct labor hours=550708/7200=$76.49(Approx)

Orders=53719/900=$59.69(Approx)

Machine hours=836316/6500=$128.66(Approx)

Hence total cost related to H9=(76.49*2400)+(59.69*200)+(128.66*3600)=$658690

Hence cost per unit=$658690/600 units

=$1097.82(Approx).